OpenStudy (awy):

Calculate y' (x)(y^4)+(x^2)(y)=(x)+(3)(y)

geerky42 (geerky42):

take derivative of both sides. once you are done, isolate y'.

OpenStudy (awy):

how exactly am i supposed to derive it

OpenStudy (awy):

i know it has something to do with the implicit derivative

geerky42 (geerky42):

let start with each terms. \(\dfrac{d}{dx} xy^4 = ?\)

geerky42 (geerky42):

yeah it is implicit derivative

OpenStudy (awy):

y^4+x*4y^3

geerky42 (geerky42):

hmm you forgot something.

geerky42 (geerky42):

remember? \(\dfrac{d}{dx}y = y'\) right? so it's supposed to be \(y^4 y' + 4xy^3 y'\)

geerky42 (geerky42):

does that make sense?

OpenStudy (awy):

hmm, why is it y' and not x' ?

geerky42 (geerky42):

because you take derivative in respect to x. so x' = d/dx x = 1

geerky42 (geerky42):

y' because we don't know what y is so we just let it y'. y is function of x.

geerky42 (geerky42):

is that clear?

OpenStudy (awy):

yes

geerky42 (geerky42):

ok now next term in left side. \(\dfrac{d}{dx} x^2y = ?\)

OpenStudy (awy):

y'*2x+y'*x^2

geerky42 (geerky42):

close. it's \( 2xy + x^2y'\)

geerky42 (geerky42):

when you take derivative of x, you dont do it to y. get it?

OpenStudy (awy):

but in x*y^4 you took the d/dx of x and got 1

geerky42 (geerky42):

\(\dfrac{d}{dx}uv = u'v + uv'\) u is x^2 and v is y. we take derivative of x^2, so we need to left y untouched. yeah d/dx of x is 1 so d/dx x*y^4 = 1*y^4 + 4 x y^3 y'

geerky42 (geerky42):

so is that clear??

OpenStudy (awy):

so i'm not quite sure when you have y'...

geerky42 (geerky42):

you get y' when you take derivative of y.

geerky42 (geerky42):

so for example: d/dx y^2 = 2y^1 y' because of chain rule; we just deal with exponent, then we deal with y itself.

OpenStudy (awy):

when we took the d/dx of x*y^4, i was supposed to use the chain rule?

geerky42 (geerky42):

and product rule. we use product rule first

geerky42 (geerky42):

for y part, we then use chain rule

OpenStudy (awy):

ok but we have to take the d/dx of x and you said to put y'? but we didnt take the d/dx of y tho

geerky42 (geerky42):

yeah, because we left y untouched. \(\dfrac{d}{dx} xy^4 = \left( \dfrac{d}{dx}x\right)y^4 + x\left(\dfrac{d}{dx}y^4\right)\) \( = 1\cdot y^4 + x\cdot 4y^3 \space y'\)

geerky42 (geerky42):

when we take derivative of y^4, we used chain rule

OpenStudy (awy):

ohh...

geerky42 (geerky42):

so you understand?

OpenStudy (awy):

yes kind of

geerky42 (geerky42):

ok try again. d/dx x^2 y

OpenStudy (awy):

so d/dx of x is 1 and d/dx of y =y' ?

geerky42 (geerky42):

yeah

geerky42 (geerky42):

we take derivative of y with respect to x. we don't know what y is equal to so we just left it y', which is another way to say d/dx y

OpenStudy (awy):

d/dx of x^2 y = (2xy) + (x^2)(y')

geerky42 (geerky42):

yeah!

OpenStudy (awy):

ok

geerky42 (geerky42):

now take derivative of right side

OpenStudy (awy):

d/dx (x) + 3y' ?

OpenStudy (awy):

i mean 1 + 3 y'

geerky42 (geerky42):

yeah, 1 + 3y'

geerky42 (geerky42):

so all together, we get \(y^4 + 4xy^3y' + 2xy + x^2y' = 1+3y'\)

geerky42 (geerky42):

now we are done with derivative. left thing for us to do is to isolate y'

geerky42 (geerky42):

can you do that?

OpenStudy (awy):

y'(4xy^3+x^2-3)+y^4+2xy-1 ?

OpenStudy (awy):

y'=(4xy^3+x^2-3)/-(y^4+2xy-) ?

OpenStudy (awy):

-(y^4+2xy+1)

geerky42 (geerky42):

yeah, now distribute negative sign in denominator, then that's your final answer.

geerky42 (geerky42):

-(y^4+2xy-1)

OpenStudy (awy):

alright i get it now, tyvm :P

geerky42 (geerky42):

no problem. glad i helped